# Multimodal Contrastive Representation Learning in Augmented Biomedical Knowledge Graphs

### v1

@@ Abstract
1 Introduction
Biomedical knowledge graphs (BKGs) are structured networks that represent intricate relationships among biological entities such as genes, proteins, diseases, and drugs (see Figure 1). Accurate link prediction within these graphs is crucial for identifying hidden relationships, discovering potential therapeutic targets, and suggesting drug repositioning opportunities Nicholson and Greene (2020); Zitnik et al. (2018); Ngo et al. (2022). These capabilities can significantly accelerate biomedical research, leading to faster clinical advancements and more effective treatments.
@@ 1 Introduction
Figure 1: The subgraph illustrates the interactions surrounding the Parathyroid hormone receptor and its connections to related drugs and diseases. Different entity types are color-coded: red nodes represent drugs, blue nodes indicate genes or proteins, and yellow nodes denote diseases. Black arrows depict drug-treatment relationships with diseases, while orange arrows represent drug-receptor interactions. This subgraph is a focused segment of a broader Biomedical Knowledge Graph, which captures the complex interconnections among various biological entities.
@@
Recent advances show that pre-trained Language Models (LMs) can act as foundational knowledge bases, storing vast amounts of factual information Petroni et al. (2019); He et al. (2024); Zhao et al. (2024); Jiang et al. (2024); Chen (2023). When used as initial embeddings, LMs provide a strong foundation for downstream tasks by incorporating pre-existing knowledge from biomedical texts and databases Wang et al. (2023). These models offer rich semantic information that can enhance graph representation learning. However, prior works on BKGs Daza et al. (2023); Lam et al. (2023) have primarily focused on using LMs for single-modality node representations, neglecting the integration of multimodal information. Moreover, while LM-derived embeddings serve as initial representations for knowledge graphs, they often lack graph topology, necessitating fine-tuning to effectively capture graph structure.
@@
In this work, we propose a novel pre-trained node representation model designed to enhance link prediction performance in Biomedical Knowledge Graphs (BKGs). Our comprehensive framework leverages the capabilities of Language Models (LMs) to generate robust entity representations, while seamlessly integrating multimodal information to enrich the contextual understanding of relationships within the graph. Specifically, we unify LM-derived embeddings for each entity and employ Graph Contrastive Learning (GCL) to optimize intra-node relationships by enhancing mutual information within individual node types. Additionally, we utilize a Knowledge Graph Embedding (KGE) model to capture inter-node information between different biological entities. A key feature of our approach is its generalizability, as the node embeddings generated by our framework encapsulate both semantic information from LMs and relational information from GCL. This dual integration ensures that embeddings maintain a rich contextual understanding, allowing the framework to generate meaningful representations even for unseen nodes, thereby facilitating more accurate link prediction for novel entities.
@@
However, our approach requires a BKG with well-defined node attributes, which are absent in most existing BKGs that lack comprehensive attributes for each entity type Chandak et al. (2023); Walsh et al. (2020). To address this limitation, we introduce PrimeKG++, an enriched knowledge graph that builds upon PrimeKG Chandak et al. (2023). PrimeKG++ enhances the original dataset by incorporating biological sequences for each entity type—amino acid sequences for proteins, nucleic acid sequences for genes, and SMILES strings for small molecules—along with comprehensive textual descriptions. This integration diversifies node attributes and improves the overall utility of the knowledge graph, providing a valuable public resource for future research in biomedical knowledge graphs.
@@
It is important to clarify that the primary focus of this paper is not on achieving state-of-the-art (SOTA) results in downstream tasks such as link prediction. Instead, we aim to propose a pretrained node representation model and demonstrate its effectiveness through comprehensive experiments. To evaluate this, we employed existing models with and without our pretrained node representations as initial inputs. Our experiments show that our pretrained node representations lead to significant performance improvements compared to random initialization or Direct LM-derived embeddings. By leveraging SOTA models for link prediction, we ensured that our comparisons were rigorous and meaningful, demonstrating the added value of our pretrained node representations within an established and high-performing framework.
@@
  • We propose a comprehensive framework that leverages pretrained Language Models (LMs) and Graph Contrastive Learning (GCL) to create robust, multimodal node embeddings for Biomedical Knowledge Graphs (BKGs).
@@
2 Related works
2.1 Knowledge Graph Embedding
@@
In the field of biomedical knowledge graph (BKGs), link prediction research aims to uncover connections among biological entities by analyzing their existing links and attributes Menon and Elkan (2011); Zitnik et al. (2018); H"ansel et al. (2023); Wang et al. (2021); Fu et al. (2021). Knowledge graph embeddings, representing entities and relations as vectors, have gained popularity for this task. While traditional models, such as ComplEx Trouillon et al. (2016) and RotatE Sun et al. (2019) have shown promising results in this link prediction task, two key constraints hinder them: firstly, they focus solely on the graph structure, ignoring valuable entity attribute information; and secondly, their reliance on predetermined embeddings for mapping entities and relations in the lookup table complicates integration with new entities. These constraints motivate us to construct a heterogeneous biomedical knowledge graph with multimodal metadata.
2.2 Biomedical Language Model
@@
In BKGs, entities can possess different modalities, such as text or biological sequences. Essentially, a molecular sequence is the exact order of smaller units (monomers) that make up a large molecule (biopolymer). Similar to a textual description, it inherently possesses a sequential relationship that Language Models (LMs) can effectively process. Recent methods rely on pre-trained language models such as BERT Devlin et al. (2019) as the backbone for the attribute encoder. Protein sequences, which are strings of amino acid letters, can be effectively processed by models like ESM-2 Lin et al. (2023) and ProteinBERT Brandes et al. (2022). For genes, which are represented by nucleotide sequences, specific language models such as Nucleotide Transformers Dalla-Torre et al. (2023) and DNABERT Ji et al. (2021) are required. Chemical structures are often represented using SMILES strings, a linear text format, which can be interpreted by models like BARTSmiles Chilingaryan et al. (2022) and MoLFormer Ross et al. (2022). For textual descriptions in the biomedical domain, models such as BioGPT Lewis et al. (2020) and BioBERT Lee et al. (2020) are leveraged to extract high semantic meaning, providing improved understanding and analysis of biomedical text. These findings inspire us to explore the potential of LMs to extract semantic information into node features in BKGs.
2.3 Graph Contrastive Learning
@@
Many Graph Neural Networks rely on supervised learning with labeled data, which is costly and labor-intensive. To address this, a few studies(e.g., DGI Veličković et al. (2018), MVGRL Hassani and Khasahmadi (2020), GMI Peng et al. (2020), and GRACE Zhu et al. (2020)) use contrastive learning techniques, introducing Graph Contrastive Learning for self-supervised graph representation learning. These methods aim to maximize mutual information between an anchor node and its semantically similar counterparts while minimizing it for dissimilar ones. In recent years, contrastive learning has gained traction in knowledge graph embedding. KGCL Yang et al. (2022) integrates knowledge graph learning with user-item interaction modeling through a joint self-supervised learning approach, improving robustness and addressing data noise and sparsity in recommendation systems. KE-GCL Zhang and Li (2022) incorporates contextual descriptions of entities and proposes adaptive sampling to refine the knowledge graph contrastive learning. MCLEA Lin et al. (2022) unifies information from various modalities and uses contrastive learning for discriminative entity representations. However, multimodal contrastive learning has not yet been explored in BKGs. In this paper, we present a novel graph representation learning framework incorporating contrastive learning for biomedical knowledge graphs.
3 PrimeKG++: An Augmented Knowledge Graph
@@
PrimeKG Chandak et al. (2023) is a multimodal knowledge graph tailored for precision medicine, encompassing over 100,000 nodes across various biological scales. It features more than 4 million relationships among these nodes, categorized into 29 distinct edge types. We chose PrimeKG for its enriched disease and drug nodes, which are augmented with clinical descriptors sourced from medical authorities. This enrichment supports our approach by providing a robust foundation for applying LM-derived embeddings, allowing for more precise and contextually relevant analyses in biomedical research. However, PrimeKG exhibits limitations, particularly in its lack of additional context or descriptive information for other biological entity types such as genes and proteins. This deficiency hinders the graph’s capacity to fully represent the complex interactions and functions inherent in these biological components.
@@
To address these limitations, we developed PrimeKG++, an enhanced version of PrimeKG that adds detailed information for three key node types: gene/protein, drug, and disease. PrimeKG++ categorizes drug data into two subtypes: molecules, represented with SMILES strings, and antibodies, identified by amino acid sequences. For the gene/protein node type, it includes protein-coding genes, annotated with amino acid sequences, and non-coding genes, represented with nucleotide sequences. Descriptions are collected for all subtypes of both drugs and genes/proteins, providing essential context. These enhancements are meticulously linked to authoritative sources such as Entrez Gene Maglott et al. (2010) at NCBI for genes/proteins and DrugBank Knox et al. (2024b) for drugs, using identifiers from PrimeKG. This detailed information allows PrimeKG++ to analyze and distinguish these entities’ complex interactions and functions with greater precision. The step-by-step production process is presented in Appendix B.
4 Method
@@
Figure 2: Overview of our proposed framework. A. Modality Embedding: Creating node attribute embeddings through domain-specific LMs. B. Contrastive Learning: Enhancement of LM-derived embeddings for specific node attributes of the same type through Fusion Module and Contrastive Learning. C. Link Prediction on KG Embedding: Utilizing the enhanced embeddings to perform link prediction tasks through a Knowledge Graph Embedding (KGE) model that learns relationships and enhances semantic information across distinct node types.
4.1 Preliminaries
In the context of knowledge graphs where entities have associated attributes across various modalities, we define a Biomedical Knowledge Graph as $G=(V,R,E,D,d)$. Here, $V$ denotes the set of nodes, collectively $\{v_{1},\dots,v_{n}\}$, with n being the number of entities, $R$ denotes the set of relations, and $E$ consists of triples $(h,r,t)$ where $h,t\in V$ and $r\in R$. The elements $h$, $r$, and $t$ represent the head, relation, and tail of a triple, respectively. $D$ represents a dataset with entities’ attributes, where each entity type has specific attributes relevant to its biological role. The partial function $d:V_{d}\rightarrow D$ maps a subset of entities $V_{d}\subseteq V$, which have available attribute data, to their respective attributes, with $d(v_{i})$ retrieving the attribute data for an entity $v_{i}$. This schema allows for tailored attribute representation, accommodating the diverse and specific data needs of different entity types within the graph. The complete list of notations is in Table 5 in Table 4.
4.2 Modality Encoding
We utilize a set of $k$ modality-specific encoders, $\{e_{1},\dots,e_{k}\}$, where each encoder $e_{i}$ corresponds to a pre-trained Language Model for a specific attribute modality $D_{i}\subseteq D$. Each encoder $e_{i}$ maps its respective attribute data into a distinct embedding space $\mathcal{X}_{i}$, formally represented as $e_{i}:D_{i}\rightarrow\mathcal{X}_{i}$. Selection of these LMs involves ensuring uniformity in embedding sizes and balancing the computational complexity with the desired level of accuracy, to optimize both integration across modalities and overall system efficiency. Specifically, we use ProtBERT Brandes et al. (2022) for protein sequences, DNABERT Ji et al. (2021) for gene sequences, MolFormer Ross et al. (2022) for molecule SMILES strings, and BioBERT Lee et al. (2020) for descriptions of all entity types. These choices ensure that the attribute embeddings leverage domain-specific knowledge encoded in the LMs, thereby enhancing the quality and applicability of the generated embeddings. During training, the LMs are frozen to reduce the number of trainable parameters. Another concern is that Knowledge Graphs are often incomplete due to undisclosed or overlooked facts. For these cases, we randomly initialize attribute embeddings.
4.3 Modality Fusing
@@
To achieve the integration of these modality-specific embeddings effectively, we utilize Attention Fusion Vaswani et al. (2023) and Relation-guided Dual Adaptive Fusion (ReDAF) Zhang et al. (2024). These fusion methods determine the contribution of each modality before combining them into a unified representation, which is essential because different modalities may carry varying levels of importance depending on the context. By assigning appropriate weights to each modality, the model can better capture the most relevant information, resulting in a more accurate and meaningful representation of the entity. Detailed descriptions of these techniques are provided in Appendix C. Regardless of the fusion method used, a simple mean operation is applied at the final stage to ensure a balanced integration of the multi-modal embeddings, allowing for a cohesive representation of each entity.
4.4 Graph Contrastive Learning
We employ Graph Contrastive Learning (GCL) models to maximize the agreement between two augmented views of the same graph, facilitating the extraction of valuable insights among nodes of identical types. We specifically explore various GCL models that are suitable for Knowledge Graphs, including Deep Graph Infomax (DGI) Veličković et al. (2018), Graph Group Discrimination (GGD) Zheng et al. (2022), and Graph Contrastive Representation Learning (GRACE) Zhu et al. (2020). Each of these models employs different strategies for training using the contrastive learning approach, as further elaborated in Appendix D. Regarding augmentation techniques, while the diffusion method has demonstrated superior effectiveness Hassani and Khasahmadi (2020), it also demands more execution time compared to alternatives. Therefore, for the sake of efficiency, we opt to mask out nodes and remove edges for quick experimentation randomly.

### v2

@@ Abstract.
Biomedical knowledge graphs, Multimodal graph representation learning, Graph contrastive learning, Medical language models, Data augmentation, Link prediction, Drug repurposing.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 24th International Workshop on Data Mining in Bioinformatics; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
BKGs are structured networks that represent intricate relationships among biological entities such as genes, proteins, diseases, and drugs (see Figure 1). Accurate link prediction within these graphs is crucial for identifying hidden relationships, discovering potential therapeutic targets, and suggesting drug repositioning opportunities (BKGs_applications, ; biolp2, ; ngo2022predicting, ). These capabilities can significantly accelerate biomedical research, leading to faster clinical advancements and more effective treatments.
@@ 1. Introduction
Figure 1. The subgraph illustrates the interactions surrounding the Parathyroid hormone receptor and its connections to related drugs and diseases. Different entity types are color-coded: red nodes represent drugs, blue nodes indicate genes or proteins, and yellow nodes denote diseases. Black arrows depict drug-treatment relationships with diseases, while orange arrows represent drug-receptor interactions. This subgraph is a focused segment of a broader Biomedical Knowledge Graph, which captures the complex interconnections among various biological entities.
@@
Recent advances show that pre-trained LMs can act as foundational knowledge bases, storing vast amounts of factual information (LMs_knowledge_bases1, ; LMs_knowledge_bases2, ; LMs_knowledge_bases4, ; LMs_knowledge_bases3, ; LMs_knowledge_bases5, ). When used as initial embeddings, LMs provide a strong foundation for downstream tasks by incorporating pre-existing knowledge from biomedical texts and databases (LMs_survey, ). These models offer rich semantic information that can enhance graph representation learning. However, prior works on BKGs (BioBLP, ; Otter-Knowledge, ) have primarily focused on using single-modality node representations for each node type (e.g., amino acid sequences for proteins, SMILES strings for drugs, and textual descriptions for diseases), overlooking the potential to integrate multiple modalities for each node type. Moreover, while LM-derived embeddings serve as initial representations for knowledge graphs, they often lack graph topology, necessitating fine-tuning to effectively capture graph structure.
@@
In this work, we propose a novel pre-trained node representation model designed to enhance link prediction performance in BKGs. Our comprehensive framework leverages the capabilities of LMs to generate robust entity representations, while seamlessly integrating multimodal information to enrich the contextual understanding of relationships within the graph. Specifically, we unify LM-derived embeddings for each entity and employ GCL to optimize intra-node relationships by enhancing mutual information within individual node types. Additionally, we utilize a KGE model to capture inter-node information between different biological entities. A key feature of our approach is its generalizability, as the node embeddings generated by our framework encapsulate both semantic information from LMs and relational information from GCL. This dual integration ensures that embeddings maintain a rich contextual understanding, allowing the framework to generate meaningful representations even for unseen nodes, thereby facilitating more accurate link prediction for novel entities.
@@
However, our approach requires a BKG with well-defined node attributes, which are absent in most existing BKGs that lack comprehensive attributes for each entity type (PrimeKG, ; Biokg, ). To address this limitation, we introduce PrimeKG++, an enriched knowledge graph that builds upon PrimeKG (PrimeKG, ). PrimeKG++ enhances the original dataset by incorporating biological sequences for each entity type—amino acid sequences for proteins, nucleic acid sequences for genes, and SMILES strings for small molecules—along with comprehensive textual descriptions. This integration diversifies node attributes and improves the overall utility of the knowledge graph, providing a valuable public resource for future research in biomedical knowledge graphs.
@@
It is important to clarify that the primary focus of this paper is not on achieving state-of-the-art results in downstream tasks such as link prediction. Instead, we aim to propose a pretrained node representation model and demonstrate its effectiveness through comprehensive experiments. To evaluate this, we employed existing models with and without our pretrained node representations as initial inputs. Our experiments show that our pretrained node representations lead to significant performance improvements compared to random initialization or Direct LM-derived embeddings. By leveraging SOTA models for link prediction, we ensured that our comparisons were rigorous and meaningful, demonstrating the added value of our pretrained node representations within an established and high-performing framework.
@@
  • We propose a comprehensive framework that leverages LMs and GCL to create robust, multimodal node embeddings for BKGs.
@@
2. Related works
2.1. Knowledge Graph Embedding
@@
In the field of BKGs, link prediction research aims to uncover connections among biological entities by analyzing their existing links and attributes (biolp1, ; biolp2, ; biolp3, ; biolp4, ; biolp5, ). Knowledge graph embeddings, representing entities and relations as vectors, have gained popularity for this task. While traditional models, such as ComplEx (ComplEx, ) and RotatE (Rotate, ) have shown promising results in this link prediction task, two key constraints hinder them: firstly, they focus solely on the graph structure, ignoring valuable entity attribute information; and secondly, their reliance on predetermined embeddings for mapping entities and relations in the lookup table complicates integration with new entities. These constraints motivate us to construct a heterogeneous biomedical knowledge graph with multimodal metadata.
2.2. Biomedical Language Model
@@
In BKGs, entities can possess different modalities, such as text or biological sequences. Essentially, a molecular sequence is the exact order of smaller units (monomers) that make up a large molecule (biopolymer). Similar to a textual description, it inherently possesses a sequential relationship that LMs can effectively process. Recent methods rely on pre-trained language models such as BERT (bert, ) as the backbone for the attribute encoder. Protein sequences, which are strings of amino acid letters, can be effectively processed by models like ESM-2 (ESM-2, ) and ProteinBERT (ProteinBERT, ). For genes, which are represented by nucleotide sequences, specific language models such as Nucleotide Transformers (Dalla-Torre2023.01.11.523679, ) and DNABERT (DNABERT, ) are required. Chemical structures are often represented using SMILES strings, a linear text format, which can be interpreted by models like BARTSmiles (BARTSmiles, ) and MoLFormer (MolFormer, ). For textual descriptions in the biomedical domain, models such as BioGPT (BioGPT, ) and BioBERT (BioBERT, ) are leveraged to extract high semantic meaning, providing improved understanding and analysis of biomedical text. These findings inspire us to explore the potential of LMs to extract semantic information into node features in BKGs.
2.3. Graph Contrastive Learning
@@
Many Graph Neural Networks rely on supervised learning with labeled data, which is costly and labor-intensive. To address this, a few studies(e.g., DGI (DGI, ), MVGRL (MVGRL, ), GMI (GMI, ), and GRACE (GRACE, )) use contrastive learning techniques, introducing Graph Contrastive Learning for self-supervised graph representation learning. These methods aim to maximize mutual information between an anchor node and its semantically similar counterparts while minimizing it for dissimilar ones. In recent years, contrastive learning has gained traction in knowledge graph embedding. KGCL (kgcl2022, ) integrates knowledge graph learning with user-item interaction modeling through a joint self-supervised learning approach, improving robustness and addressing data noise and sparsity in recommendation systems. KE-GCL (KEGCL, ) incorporates contextual descriptions of entities and proposes adaptive sampling to refine the knowledge graph contrastive learning. MCLEA (multimodal_gcl, ) unifies information from various modalities and uses contrastive learning for discriminative entity representations. However, multimodal contrastive learning has not yet been explored in BKGs. In this paper, we present a novel graph representation learning framework incorporating contrastive learning for biomedical knowledge graphs.
3. PrimeKG++: An Augmented Knowledge Graph
@@
PrimeKG (PrimeKG, ) is a multimodal knowledge graph tailored for precision medicine, encompassing over 100,000 nodes across various biological scales. It features more than 4 million relationships among these nodes, categorized into 29 distinct edge types. We chose PrimeKG for its enriched disease and drug nodes, which are augmented with clinical descriptors sourced from medical authorities. This enrichment supports our approach by providing a robust foundation for applying LM-derived embeddings, allowing for more precise and contextually relevant analyses in biomedical research. However, PrimeKG exhibits limitations, particularly in its lack of additional context or descriptive information for other biological entity types such as genes and proteins. This deficiency hinders the graph’s capacity to fully represent the complex interactions and functions inherent in these biological components.
@@
To address these limitations, we developed PrimeKG++, an enhanced version of PrimeKG that adds detailed information for three key node types: gene/protein, drug, and disease. PrimeKG++ categorizes drug data into two subtypes: molecules, represented with SMILES strings, and antibodies, identified by amino acid sequences. For the gene/protein node type, it includes protein-coding genes, annotated with amino acid sequences, and non-coding genes, represented with nucleotide sequences. Descriptions are collected for all subtypes of both drugs and genes/proteins, providing essential context. These enhancements are meticulously linked to authoritative sources such as Entrez Gene (NCBI, ) at NCBI for genes/proteins and DrugBank (DrugBank, ) for drugs, using identifiers from PrimeKG. This detailed information allows PrimeKG++ to analyze and distinguish these entities’ complex interactions and functions with greater precision.
4. Method
@@
Figure 2. Overview of our proposed framework. A. Modality Embedding: Creating node attribute embeddings through domain-specific LMs. B. Contrastive Learning: Enhancement of LM-derived embeddings for specific node attributes of the same type through Fusion Module and Contrastive Learning. C. Link Prediction on KG Embedding: Utilizing the enhanced embeddings to perform link prediction tasks through a Knowledge Graph Embedding (KGE) model that learns relationships and enhances semantic information across distinct node types.
4.1. Preliminaries
In the context of knowledge graphs where entities have associated attributes across various modalities, we define a Biomedical Knowledge Graph as $G=(V,R,E,D,d)$. Here, $V$ denotes the set of nodes, collectively $\{v_{1},\dots,v_{n}\}$, with n being the number of entities, $R$ denotes the set of relations, and $E$ consists of triples $(h,r,t)$ where $h,t\in V$ and $r\in R$. The elements $h$, $r$, and $t$ represent the head, relation, and tail of a triple, respectively. $D$ represents a dataset with entities’ attributes, where each entity type has specific attributes relevant to its biological role. The partial function $d:V_{d}\rightarrow D$ maps a subset of entities $V_{d}\subseteq V$, which have available attribute data, to their respective attributes, with $d(v_{i})$ retrieving the attribute data for an entity $v_{i}$. This schema allows for tailored attribute representation, accommodating the diverse and specific data needs of different entity types within the graph.
4.2. Modality Encoding
We utilize a set of $k$ modality-specific encoders, $\{e_{1},\dots,e_{k}\}$, where each encoder $e_{i}$ corresponds to a pre-trained Language Model for a specific attribute modality $D_{i}\subseteq D$. Each encoder $e_{i}$ maps its respective attribute data into a distinct embedding space $\mathcal{X}_{i}$, formally represented as $e_{i}:D_{i}\rightarrow\mathcal{X}_{i}$. Selection of these LMs involves ensuring uniformity in embedding sizes and balancing the computational complexity with the desired level of accuracy, to optimize both integration across modalities and overall system efficiency. Specifically, we use ProtBERT (ProteinBERT, ) for protein sequences, DNABERT (DNABERT, ) for gene sequences, MolFormer (MolFormer, ) for molecule SMILES strings, and BioBERT (BioBERT, ) for descriptions of all entity types. These choices ensure that the attribute embeddings leverage domain-specific knowledge encoded in the LMs, thereby enhancing the quality and applicability of the generated embeddings. During training, the LMs are frozen to reduce the number of trainable parameters. Another concern is that Knowledge Graphs are often incomplete due to undisclosed or overlooked facts. For these cases, we randomly initialize attribute embeddings.
4.3. Modality Fusing
@@
To achieve the integration of these modality-specific embeddings effectively, we utilize Attention Fusion (Attention, ) and Relation-guided Dual Adaptive Fusion (ReDAF) (ReDAF, ). These fusion methods determine the contribution of each modality before combining them into a unified representation, which is essential because different modalities may carry varying levels of importance depending on the context. By assigning appropriate weights to each modality, the model can better capture the most relevant information, resulting in a more accurate and meaningful representation of the entity. Regardless of the fusion method used, a simple mean operation is applied at the final stage to ensure a balanced integration of the multi-modal embeddings, allowing for a cohesive representation of each entity.
4.4. Graph Contrastive Learning
We employ GCL models to maximize the agreement between two augmented views of the same graph, facilitating the extraction of valuable insights among nodes of identical types. We specifically explore various GCL models that are suitable for Knowledge Graphs, including Deep Graph Infomax (DGI) (DGI, ), Graph Group Discrimination (GGD) (GGD, ), and Graph Contrastive Representation Learning (GRACE) (GRACE, ). Each of these models employs different strategies for training using the contrastive learning approach. Regarding augmentation techniques, while the diffusion method has demonstrated superior effectiveness (MVGRL, ), it also demands more execution time compared to alternatives. Therefore, for the sake of efficiency, we opt to mask out nodes and remove edges for quick experimentation randomly.